# Knowledge Independence Breeds Disruption but Limits Recognition

### v1

@@ Acknowledgments
Figure S11: Papers with higher KI are underrepresented among cited references and receiving delayed citations. (A-F) For papers published in the year $T$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline.
Figure S12: The distribution and evolution pattern of knowledge independence in SciSciNet dataset. (A) The KI distribution of 49,579,961 papers with at least two references published between 1950 and 2021 in the SciSciNet. (B) The evolving downtrend of KI over time across disciplines (the only exception is observed for Art). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S13: Knowledge independence is associated with scientific disruption in SciSciNet dataset. (A) For the 42,468,785 papers in the SciSciNet that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. (B) The association between KI and disruption persists, regardless of the impact of the focal paper. Here, impact is measured as the number of citations received within the first five years after publication, denoted by $C_{5}$. (C) The association between KI and disruption persists across decades. (D) The association between KI and disruption persists, regardless of discipline. Bootstrapped 95% confidence intervals are shown as shaded bands in panels A-D. (E) The ATT (average treatment effect on the treated) matrix of KI on disruption via propensity score matching. Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Figure S14: Team’s preference for knowledge independence in SciSciNet dataset. (A-C) The KI values in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative KI percentile decreases with team size (panel A) and team distance (panel B), and increases with team freshness (panel C). Here we calculate the geographic team distance by the coordinate information of affiliations, and categorize teams into two types—onsite ($\leq 100$ km) and remote ($>100$ km) teams (panel B). The relationship between KI and average collaboration distance is displayed in inset of panel B. (D-F) Solid lines depict the relationships between disruption and team properties with fixed KI values, alongside the uncontrolled cases (insets). Dashed lines represent linear fitted curves. Teams with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are flatter compared to the uncontrolled curves (insets). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S15: Knowledge independence explains the negative relationship between disruption and impact in SciSciNet dataset. We observe a universal negative relationship between papers’ disruption and impact, regardless of the journal rank (panel A), time period (panel B), and discipline (panel C). Here, a paper’s impact is measured by the percentile of its total citations within the same publication year and discipline, and the journal rank is based on Scimago Journal Rank. (D) For papers published in the year $T=2000$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline (see figure S21 for different values of $T$). (E) The citation percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative citation percentile decreases with KI percentile. (F) A paper with a higher KI is more likely to become a Sleeping Beauty (SB coefficient, blue curve, left y-axis) and takes a longer silence to usher in the citation burst (Awaking time, green curve, right y-axis). (G) Solid lines depict the relationships between disruption and impact with fixed KI values, alongside the uncontrolled case (inset). Dashed lines represent linear fitted curves. For papers of comparable impact, those with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are significantly flatter compared to the uncontrolled curve (inset). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S16: Distribution of reference list length in SciSciNet dataset. The distribution of reference count for articles recorded in SciSciNet follows a stretched exponential pattern, with a large number of papers containing relatively short reference lists.
Figure S17: The association between disruption and KI is robust across different reference list lengths in SciSciNet dataset. (A-B) When controlling for the percentile of reference count, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference count. (C-D) When controlling for the reference count within the bottom group ($<=10$), where each reference count induces a limited range of KI values, both the percentile and positive ratio of disruption continue to increase with KI across all groups of reference count.
Figure S18: The association between disruption and KI is robust across alternative measures in SciSciNet dataset. (A-D) Upper panel: from the perspective of percentile, all disruption measures show a consistent increasing trend with all KI measures. Lower panel: from the perspective of positive ratio, all disruption measures similarly increase with all KI measures. The positive ratio represents the likelihood of having a higher proportion of type $i$ references compared to type $j$ references. Consequently, the variation curves for $D_{1}$ and $D_{3}$ align with those for $D_{0}$ and $D_{2}$, respectively.
Figure S19: The association between disruption and KI is robust when controlling for focal paper’s novelty and reference’s properties in SciSciNet dataset. (A-D) When controlling for focal paper’s novelty (panel A), reference age (panel B), reference impact $C_{5}$ (panel C), and reference disruption (panel D), respectively, both the percentile (blue curves) and positive ratio (green curves) of disruption continue to increase with KI across all levels of above covariates.
Figure S20: The association between disruption and KI is robust when controlling for team compositions in SciSciNet dataset. (A-C) When controlling for team size (panel A), geographic distance (panel B), and collaboration freshness (panel C) separately, both the percentile (blue curves) and positive ratio (green curves) of disruption continue to increase with KI across all levels of above team compositions. Additionally, papers produced by small, onsite, and fresh teams tend to exhibit slightly higher disruption.
Figure S21: Papers with higher KI are underrepresented among cited references and receiving delayed citations in SciSciNet dataset. (A-G) For papers published in the year $T$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline.
Figure S22: Knowledge independence is awarded but underfunded. (A) Based on the dataset for Nobel laureates  (?), We match 425 Nobel Prize-winning papers that cover the time period 1902–2016 in SciSciNet dataset, including 78 papers in Physics, 133 papers in Chemistry, and 214 papers in Medicine. The probability of winning the Nobel prize increases with the KI value of papers. (B) Based on the funding information recorded in SciSciNet dataset, we extract 263,5812 papers funded by NIH and 92,9177 papers funded by NSF. Here we repeat counting when a paper is funded by both agencies. The probability of being funded decreases with the KI value of papers.
Table S1: Normalized mediation regression on reference disruption. We perform normalized mediation analysis to assess the effect of reference disruption on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by reference disruption, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Reference disruption KI 0.2086*** 0.2318*** 0.4404***
@@
Reference disruption Novelty -0.0008*** 0.4412*** 0.4404***
@@
Reference disruption Reference count 0.0227*** 0.4177*** 0.4404***
@@
Reference disruption Reference age -0.0023*** 0.4427*** 0.4404***
@@
Reference disruption Reference $C_{5}$ 0.0005*** 0.4399 0.4404***
@@
Reference disruption Team size 0.0001*** 0.4403*** 0.4404***
@@
Reference disruption Team distance 0.0018*** 0.4386*** 0.4404***
@@
Reference disruption Team freshness -0.0000*** 0.4404*** 0.4404***
@@
Table S2: Normalized mediation regression on team size. We perform normalized mediation analysis to assess the effect of team size on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team size KI -0.0043*** -$0**$0123***
@@
Team size Novelty -$0**$0123*** -0.0123***
@@
Team size Reference count -0.0002*** -0.0121*** -0.0123***
@@
Team size Reference age -0.0018*** -0.0105*** -0.0123***
@@
Team size Reference $C_{5}$0.0004*** -0.0119 -0.0123***
@@
Team size Team distance -0.0037*** -0.0087*** -0.0123***
@@
Team size Team freshness -$0**$0123*** -0.0123***
@@
Table S3: Normalized mediation regression on team distance. We perform normalized mediation analysis to assess the effect of team distance on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team distance KI -0.0255*** -0.0318*** -0.0573***
@@
Team distance Novelty -$0**$0573*** -0.0573***
@@
Team distance Reference count -0.0084*** -0.0489*** -0.0573***
@@
Team distance Reference age -0.0027*** -0.0547*** -0.0573***
@@
Team distance Reference $C_{5}$0.0003*** -0.0571*** -0.0573***
@@
Team distance Team size -0.0007*** -0.0566*** -0.0573***

### v2

@@ Acknowledgements
Figure 12: $|$ Papers with higher KI are underrepresented among cited references and receiving delayed citations. (A-H) For papers published in the year $T$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline.
Table 1: $|$ Normalized mediation regression on team size.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team size KI -0.0196*** -0.0149*** -0.0345*** 0.5680
@@
Team size Novelty -0.0000*** -0.0345*** -0.0345*** 0.0000
@@
Team size Interdisciplinarity -0.0020*** -0.0325*** -0.0345*** 0.0575
@@
Team size Reference count -0.0049*** -0.0296*** -0.0345*** 0.1413
@@
Team size Reference age -0.0011*** -0.0333*** -0.0345*** 0.0330
@@
Team size Team distance -0.0092*** -0.0253*** -0.0345*** 0.2678
@@
Team size Team freshness 0.0082*** -0.0426*** -0.0345*** 0.2368
@@
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team size on disruption. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table 2: $|$ Normalized mediation regression on team distance.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team distance KI -0.0330*** -0.0310*** -0.0640*** 0.5151
@@
Team distance Novelty 0.0000*** -0.0640*** -0.0640*** 0.0000
@@
Team distance Interdisciplinarity 0.0008*** -0.0648*** -0.0640*** 0.0121
@@
Team distance Reference count -0.0181*** -0.0459*** -0.0640*** 0.2824
@@
Team distance Reference age -0.0002*** -0.0638*** -0.0640*** 0.0036
@@
Team distance Team size -0.0039*** -0.0601*** -0.0640*** 0.0605
@@
Team distance Team freshness 0.0005*** -0.0645*** -0.0640*** 0.0079
@@
  • $\dagger$
    We perform normalized mediation analysis to assess the effect of team distance on disruption. The candidate mediating variables include: KI, novelty, interdisciplinarity, reference count, reference age, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Table 3: $|$ Normalized mediation regression on team freshness.
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$) ||ACME / TE||
Team freshness KI 0.0465*** 0.0480*** 0.0946*** 0.4922
@@
Team freshness Novelty 0.0000*** 0.0946*** 0.0946*** 0.0000
@@
Team freshness Interdisciplinarity 0.0001*** 0.0945*** 0.0946*** 0.0013
@@
Team freshness Reference count 0.0136*** 0.0810*** 0.0946*** 0.1439
@@
Team freshness Reference age 0.0000*** 0.0946*** 0.0946*** 0.0002
@@
Team freshness Team size -0.0035*** 0.0981*** 0.0946*** 0.0375
@@
Team freshness Team distance -0.0003*** 0.0949*** 0.0005*** 0.0036